# SSF-Net: Spatial-Spectral Fusion Network with Spectral Angle Awareness for Hyperspectral Object Tracking

### v1

@@ IV-B Comparison with State-of-the-arts
The experimental results are listed in Table I. With no consideration of HS feature extraction, the performance of BAE-Net using only spectral attention is 0.606 of the AUC score and 0.878 of DP_20. The tracking performance of SST-net, a band regrouping method that considers spatial, spectral, and temporal information simultaneously, can achieve the 0.623 AUC score and 0.917 DP_20, which compensates for the lack of single spectral attention. SEE-Net regroups bands from the perspective of autoencoder, further achieving the performance of 0.654 of AUC score and 0.907 of DP_20. It can be seen that using only the channel attention mechanism or autoencoder-like structures has limited ability in mining band relationships, which yields the general tracking performance. Methods based on band regrouping are not sufficient to extract robust HS features and using existing RGB trackers to infer the modeling of the spectral relationships will limit the performance of HS tracking. By jointly considering the spectral attention mechanism and autoencoder structure, SPIRIT provides more effective tracking results with an AUC score of 0.679 and DP_20 of 0.925. The perception of spectral relationships and the Transformer-based prediction head further improve the performance. Our method not only considers the fusion of spectral and spatial information, but also designs a spectral sensing module for tracking and prediction, and the experimental results achieve the best, with the AUC score of 0.680 and DP_20 of 0.939. This proves the effectiveness of SSF-Net, where $S^{2}$ and SAFM can extract more discriminative features, while SAAM can more accurately predict the position of objects.
@@
Comparison with RGB Trackers. In addition, we also compare the SSF-Net with advanced RGB trackers, including SiamGAT, OSTrack, SimTrack, and TransT. The false-color images of the HOTC dataset are used to perform experiments. As shown in Table I, the SSF-Net provides better performance compared with the RGB tracker. Due to the difficulty in handling factors such as lighting changes and background interference, the RGB trackers are generally weaker than HS trackers in performance, indicating the advantages of HS data in processing challenging scenes.
@@
Figure 8: The feature distribution visualization of SAAL. The red points denote the TI samples and the blue and gray points denote the positive and negative SI samples.s
Figure 9: Tracking results on the false-color images of playground, forest, car3, and coke scenarios.
Figure 10: Attribute-based comparison of success plots. (a) Success plot of background clutters (BC). (b) Success plot of deformation (DEF). (c) Success plot of fast motion (FM). (d) Success plot of illustration Variation (IV). (e) Success plot of in-plane rotation (IPR). (f) Success plot of low resolution (LR). (g) Success plot of motion blur (MB). (h) Success plot of occlusion (OCC). (i) Success plot of out-of-plane clutters (OPR). (j) Success plot of out-of-view (OV). (k) Success plot of scale variation (SV).
@@
1) Heatmap Visualization. To demonstrate the effectiveness of our proposed SAAM intuitively, we utilize the Grad-CAM++ [47] for feature visualization of classification heads and SAAM. The brighter the color, the higher the level of attention received by the network, which reflects the contribution of tracking performance in this region. The four scenarios are selected including car2, toy2, playground and rider1. The experimental results are shown in Figure 7. The first and fourth columns are the original images. The ground-truth bounding box is framed by red lines. The second and fifth columns are SAA maps and the third and sixth columns are CLS maps. The brighter the color, the higher attention received by the network, reflecting the contribution of tracking performance in this region. Due to the visually similar cars and people in the car2 and playground scenes, traditional CLS prediction heads may be confused with tracking objects, as shown in the CLS map of car2 in the third line and the CLS map of playground in the second line. The prediction head also responds to similar objects that appear, which may be detrimental to improving tracking performance. The SAAM focuses on spectral similarity calculation, which can handle the problem of similar appearances of different objects. We can see that although SAAM also responds to other similar objects, such as the SAA map of car2 in the third row and the SAA map of playground in the second row, the area with the highest response is still the tracking object, reflecting the robustness of the SAAM algorithm in handling similar objects. In addition, it can be seen from toy2 and rider1 scenarios that SAAM has a large receptive field and can perceive complete objects, avoiding the local overfitting of objects.
@@
2) Feature Distribution Comparison. To prove the validity of our proposed SAAL, we also select frames from three different clips and use the t-SNE methods to visualize the spectral features of TI and different region features of SI. The results are shown in Figure 8. The blue box in SI represents the area to be detected where the tracked object appears, while the red box in TI represents the object to be tracked, corresponding to blue and red points in the feature space, respectively. Local features that do not belong to the area to be detected are represented by gray points. After training with SAAL, the TI and SI samples can come closer in the feature space. With the proposed SAAL, the tracker can achieve better similarity modeling ability and correctly detect the object to be tracked.
@@
3) Tracking Results. We select videos from four different scenarios, including playground, forest, car3, and coke to show the visualization results of advanced RGB and HS trackers. The result is shown in Figure 9. The ground-truth bounding box is framed by black lines. The subsequent tracking results are obtained from the first frames based on the positions of labels, which are shown from the second column to the fourth column. Firstly, for playground scenes where occlusion is the main challenge, both the HS tracker and RGB tracker are susceptible to visual interference from similar objects, for example, frames 181 and 188, resulting in incorrect objects. Our proposed SSF-Net, due to its focus on the similarity between bands, has robustness in processing visually similar objects and can effectively solve the occlusion problem of similar objects. Then, we can also see from the scene of car3 that most trackers struggle to handle the scale changes of objects, especially for tracking small objects. SSF-Net can extract more spectral information, ensuring feature integrity in the event of visual information loss, thereby maintaining tracking performance. In addition, the results of fruit and coke scenes also show the effectiveness of SSF-Net in dealing with object rotation and background clutter.
@@
Attribute-Based Comparison. To demonstrate the ability to handle the challenging factors of SSF-Net, Figure 10 shows the success plot results of 11 attribute-based experiments. It can be observed that the SSF-Net achieved the top two results in 10 out of 11 scenarios. Among them, SSF-Net performs significantly better than other trackers in scenarios where spatial information is disrupted, such as BC, SV, FM, LR, DEF, and MB, which shows the robustness of the algorithm equipped with SAAM against visual interference. Due to the fusion of spectral and visual information, the SSF-Net can respond to objects in a transient occlusion state and therefore has superior performance in scenarios such as OCC and IV. For the two rotating scenes of IPR and OPR, although the visual features of the object have changed, the HS tracker can extract spectral features. In particular, SSF-Net captures the object based on spectral similarity, thus having better performance. However, the performance of SSF-Net in OV is poor because the algorithm is difficult to propose robust features for objects beyond the receptive field, and cannot integrate spectral and spatial information. Nevertheless, SSF-Net still achieves an overall performance of 0.680 AUC score and 0.939 DP_20, which demonstrates the effectiveness of the proposed methods.
IV-C Ablation Experiments
In this subsection, to evaluate the effectiveness of our methods, several ablation experiments are conducted.
@@ IV-C Ablation Experiments
Evaluation of Each Component. Firstly, we reproduce the bi-stream ResNet-50 network proposed by Liu et al. [12] as the baseline, which is marked as “B”. Then, we replace the ResNet-50 network of HS stream with $S^{2}$, which is annotated as “S”, and conduct experiments under the same fusion module SSATTN [12]. The experimental results are listed in Table II. Compared with the baseline model, $S^{2}$ improves the AUC score by 0.012 and DP_20 by 0.008. Then, based on “S”, we replace SSATTN with SAFM to demonstrate the superiority of SAFM in intra-modality and cross-modality spectral attention, which is denoted as “F”. The AUC score yields 0.675 and DP_20 yields 0.934, demonstrating the effectiveness of SAFM. Furthermore, The SAAM is incorporated into “F”, which is annotated as “A”, and the SAAL is used for training. The experimental results have further improved, yielding an AUC score of 0.680 and DP_20 of 0.939. It demonstrates the effectiveness of $S^{2}$ in extracting spectral information, the superiority of SAFM over intra-modality attention, and the ability of SAAM to supplement the shortcomings of existing tracking heads in object localization prediction.
@@
TABLE II: ablation experimental results with the baseline model. $\Delta$ indicates the amount of variation.
@@
TABLE III: ablation experimental results with the baseline model. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
TABLE IV: ablation experimental results of different fusion methods. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
TABLE V: ablation experimental results of different prediction strategy. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
Evaluation of Different Backbone Settings. To assess the impact of varying backbone architectures, we conduct experiments with the individual spatial and spectral branches of the $S^{2}$, and removing the residual components. The experimental results are listed in Table III. We first verify the effectiveness of the residual structure, and the experimental results achieve 0.673 of the AUC score and 0.932 of DP_20, which shows the importance of residual structure. Next, we explore the branches independently for feature extraction. The spatial branch, solely based on depthwise separable convolutions, lacks comprehensive channel information, resulting in a lower AUC score and DP_20 of 0.470 and 0.786, respectively. Meanwhile, the spectral branch alone achieved an AUC score of 0.658 and a DP_20 of 0.913, illustrating that while effective, the performance of each branch is not as robust as when combined. When conducting experiments with the complete $S^{2}$ structure, the model performance achieved the AUC score of 0.675 and DP_20 of 0.934, respectively. It demonstrates that the integrated approach of $S^{2}$ to capturing both spatial and spectral information offers a more comprehensive feature representation, thereby confirming the effectiveness of the proposed $S^{2}$.
@@
Evaluation of Different Fusion Methods. The feature fusion is a key aspect of cross-modality tracking. To show the effectiveness of our proposed SAFM, we conduct experiments with various fusion methods, including element-wise addition (ADD), concatenation (CONCAT), and state-of-the-art spectral attention methods including SE-Net [48], SaE-Net [49], and ECA-Net [35]. The results of these experiments are presented in Table IV. The simplest methods, addition, and concatenation, without any attention mechanisms, exhibit the least impressive performance. The current spectral attention mechanisms focus on enhancing within-modality features and fall short in harnessing cross-modality feature interactions. Consequently, the most effective result among them is achieved by ECA-Net, yielding the AUC score and DP_20 of 0.673 and 0.932, respectively. However, these scores are still 0.2% lower than those achieved by SAFM. Such a comparison shows the advantages of SAFM in the context of feature fusion, confirming that our module not only refines intra-modality representation but also effectively captures and leverages the interrelations across modalities.
@@
Evaluation of Different Prediction Strategy. In this section, we evaluate the performance of our weighted prediction approach. Drawing upon the “S” model configuration, we compare four different prediction strategies: using the HS branch alone, using the RGB branch alone, averaging the predictions from both, and applying a weighted prediction method. Experimental results are listed in Table V. It can be seen that using the HS branch to predict only achieves the AUC score of 0.230 and DP_20 of 0.510. This was expected, as the $S^{2}$ is initially trained from scratch on the HOTC dataset, making it challenging for the HS branch alone to develop robust feature extraction capabilities. Then, the result of the RGB branch is 0.623 of the AUC score and DP_20 of 0.871. Due to the parameter freezing of the RGB branch, only the knowledge of the RGB domain is used for false-color image tracking, resulting in poor tracking performance. For average prediction, although the prediction results of both modalities are comprehensively considered, the tracking performance reached an AUC score of 0.639 and DP_20 of 0.915. However, in different scenarios, the contributions of HS and RGB modality to the results should be different. So we used the weighted prediction method to adjust the weights of the two modality predictions adaptively, and the final results are the AUC score of 0.680 and DP_20 of 0.939. This further shows the effectiveness of the weighted prediction method.

### v2

@@ IV-B Comparison with State-of-the-arts
The experimental results are listed in Table I. Without consideration of HS feature extraction, the performance of BAE-Net using only spectral attention is 0.606 of the AUC score and 0.878 of DP_20. The tracking performance of SST-net, a band regrouping method that considers spatial, spectral, and temporal information simultaneously, can achieve 0.623 of AUC score and 0.917 of DP_20, which compensates for the lack of spectral attention. SEE-Net regroups bands from the perspective of an autoencoder, further achieving 0.654 of the AUC score and 0.907 of DP_20. It can be seen that using only the channel attention mechanism or autoencoder-like structures has limited ability in mining band relationships, which yields the general tracking performance. Band grouping-based methods are not powerful in extracting robust HS features, and using existing RGB trackers to infer the modeling of the spectral relationships limits the performance of HS tracking. By jointly considering the spectral attention mechanism and autoencoder structure, SPIRIT provides more effective tracking results with an AUC score of 0.679 and DP_20 of 0.925. The perception of spectral relationships and the Transformer-based prediction head further improve the performance. Our method not only considers the fusion of spectral and spatial information but also designs a spectral sensing module for tracking and prediction, and the experimental results achieve the best, with the AUC score of 0.680 and DP_20 of 0.939. This shows the effectiveness of SSF-Net, where $S^{2}$ and SAFM can extract more discriminative features, while SAAM can more accurately predict the position of objects.
Besides, to prove the effectiveness of our methods further, we conduct comparative experiments on the HOTC-2024-RedNIR and BihoT datasets. The experimental results are listed in Table II. Our method outperforms most existing methods, achieving an AUC of 0.357 and a DP_20 of 0.463 on the HOTC-2024-RedNIR dataset, and our method surpasses other methods in terms of AUC and DP_20 on the BihoT dataset. This improvement is attributed to our proposed modules ($S^{2}$, SAFM, and SAAM), which enable the model to form joint spatial-spectral features of the objects. Additionally, for camouflaged objects with greater tracking difficulty, our method leverages SAAM to perceive spectral similar regions between the template and the search features, further enhancing tracking performance, which demonstrates the effectiveness and generalizability of our proposed methods.
Figure 8: The feature distribution visualization of SAAL. The red points denote the TI samples, and the blue and gray points denote the positive and negative SI samples.
@@
Comparison with RGB Trackers. In addition, we also compare the SSF-Net with advanced RGB trackers, including SiamGAT, OSTrack, SimTrack, and TransT. The false-color images of the HOTC-2020 dataset are used to perform experiments. As shown in Table I, the SSF-Net provides better performance compared with the RGB tracker. Due to the difficulty in handling factors such as lighting changes and background interference, the RGB trackers are generally weaker than HS trackers in processing challenging scenes.
@@
1) Heatmap Visualization. To demonstrate the effectiveness of our proposed SAAM intuitively, we utilize the Grad-CAM++ [57] for feature visualization of classification heads and SAAM. The brighter the color, the higher the level of attention received by the network, which reflects the contribution of tracking performance in this region. The four scenarios are selected including car2, toy2, playground and rider1. The experimental results are shown in Figure 7. The first and fourth columns are the original images. The ground-truth bounding box is framed by red lines. The second and fifth columns are SAA maps, and the third and sixth columns are CLS maps. The brighter the color, the higher the attention received by the network, reflecting the contribution of tracking performance in this region. Due to the visually similar cars and people in the car2 and playground scenes, traditional CLS prediction heads may be confused with tracking objects, as shown in the CLS map of car2 in the third line and the CLS map of playground in the second line. The prediction head also responds to similar objects that appear, which may be detrimental to improving tracking performance. The SAAM focuses on spectral similarity calculation, which can handle the problem of similar appearances of different objects. We can see that although SAAM also responds to other similar objects, such as the SAA map of car2 in the third row and the SAA map of playground in the second row, the area with the highest response is still the tracking object, reflecting the robustness of the SAAM algorithm in handling similar objects. In addition, it can be seen from toy2 and rider1 scenarios that SAAM has a large receptive field and can perceive entire objects, avoiding the local overfitting of objects.
Figure 9: Tracking results on the HOTC-2020 datasets of the false-color images under the playground, forest, car3, and coke scenarios.
Figure 10: Failure cases on the HOTC-2024-RedNIR datasets of the false-color images under the football1, oranges5, and rainystreet12 scenarios.
Figure 11: Attribute-based comparison of success plots. (a) Success plot of background clutter (BC). (b) Success plot of deformation (DEF). (c) Success plot of fast motion (FM). (d) Success plot of illustration Variation (IV). (e) Success plot of in-plane rotation (IPR). (f) Success plot of low resolution (LR). (g) Success plot of motion blur (MB). (h) Success plot of occlusion (OCC). (i) Success plot of out-of-plane clutters (OPR). (j) Success plot of out-of-view (OV). (k) Success plot of scale variation (SV).
@@
2) Feature Distribution Comparison. To prove the validity of our proposed SAAL, we also select frames from three different clips and use the t-SNE methods to visualize the spectral features of TI and different region features of SI. As shown in Figure 8, the blue box in SI represents the area to be detected where the tracked object appears, while the red box in TI represents the object to be tracked, corresponding to blue and red points in the feature space, respectively. Local features that do not belong to the area to be detected are represented by gray points. After training with SAAL, the TI and SI samples can come closer in the feature space. With the proposed SAAL, the tracker can achieve better similarity modeling ability and correctly detect the object to be tracked.
@@
3) Tracking Results. We select videos from four different scenarios, including playground, forest, car3, and coke, to show the visualization results of advanced RGB and HS trackers. As shown in Figure 9, the ground-truth bounding box is framed by black lines. The subsequent tracking results are obtained from the first frames based on the positions of labels, which are shown from the second to the fourth columns. Firstly, for playground scenes where occlusion is the main challenge, both the HS tracker and RGB tracker are susceptible to visual interference from similar objects, for example, frames 181 and 188, resulting in incorrect objects. Our proposed SSF-Net, due to its focus on the similarity between bands, is robust in processing visually similar objects and can effectively solve the occlusion problem of similar objects. Then, we can also see from the scene of car3 that most trackers struggle to handle the scale changes of objects, especially for tracking small objects. SSF-Net can extract more spectral information, ensuring feature integrity in the event of visual information loss, thereby maintaining tracking performance. In addition, the results of fruit and coke scenes also show the effectiveness of SSF-Net in dealing with object rotation and background clutter. Besides, we also visualize the failure cases on the HOTC-2024-RedNIR dataset, which are shown in Figure 10, When an object moves rapidly or has a small size (as shown in the first row), the useful spectral and spatial features tend to be compressed, reducing the discriminability of the extracted features. This degradation negatively impacts the tracking performance of the proposed method. Additionally, when tracking multiple similar objects (as shown in the second row), the similarity in spatial and spectral information may lead to ambiguous tracking clues, making it difficult to distinguish between objects accurately. This can result in tracking failure, especially in scenarios where objects exhibit high spectral resemblance.
@@
Attribute-based Comparison. To demonstrate the ability to handle the challenging factors of SSF-Net, Figure 11 shows the success plot results of 11 attribute-based experiments. It can be observed that the SSF-Net achieved the top two results in 10 out of 11 scenarios. Among them, SSF-Net significantly outperforms other trackers in scenarios where spatial information is disrupted, such as BC, SV, FM, LR, DEF, and MB, which shows the robustness of the algorithm equipped with SAAM against visual interference. Due to the fusion of spectral and visual information, the SSF-Net can respond to objects in a transient occlusion state and has superior performance in scenarios such as OCC and IV. For the two rotating scenes of IPR and OPR, although the visual features of the object have changed, the HS tracker can extract spectral features. In particular, SSF-Net captures the object based on spectral similarity, thus yielding better performance. However, the performance of SSF-Net in OV is poor because the algorithm makes it difficult to extract robust features for objects beyond the receptive field and struggles to integrate spectral and spatial information. Nevertheless, SSF-Net still achieves an overall performance of 0.680 AUC score and 0.939 DP_20, which demonstrates the effectiveness of the proposed methods.
IV-C Running Cost Comparison.
To demonstrate the efficiency of our proposed method, we conduct running cost experiments on the HOTC-2024-RedNIR dataset. The experimental results using FLOPs and frames per second (FPS) as the evaluation indicators are reported in Table II. All the trackers are run on a Windows machine with an NVIDIA 4090 GPU. We can see that the band regrouping-based methods (e.g., SEE-Net and SENSE) rely on the simultaneous inference of multiple false-color images, resulting in higher computational and parameter complexities compared to our proposed method, ultimately leading to lower FPS. Additionally, MMF-Net, which integrates HSI, RGB, and material information, requires extensive preprocessing steps such as spectral unmixing, significantly slowing down its tracking speed. In contrast, SSF-Net achieves high tracking performance while maintaining efficient processing, demonstrating the effectiveness and efficiency of the proposed method.
IV-D Ablation Experiments
@@
Evaluation of Each Component. To evaluate the effectiveness of the proposed component, we conduct ablation experiments on two datasets: HOTC-2020 and HOTC-2024-RedNIR. Firstly, we reproduce the bi-stream ResNet-50 network proposed by Liu et al. [14] as the baseline, which is marked as “B”. Then, we replace the ResNet-50 network of HS stream with $S^{2}$, which is annotated as “S”, and conduct experiments under the same fusion module SSATTN [14]. The experimental results are listed in Table III and Table IV. As for HOTC-2020, compared with the baseline model, $S^{2}$ improves the AUC score by 0.012 and DP_20 by 0.008. Then, based on “S”, we replace SSATTN with SAFM to demonstrate the superiority of SAFM in intra-modality and cross-modality spectral attention, which is denoted as “F”. The AUC score yields 0.675, and DP_20 yields 0.934, demonstrating the effectiveness of SAFM. Furthermore, the SAAM is incorporated into “F”, which is annotated as “A”, and the SAAL is used for training. The experimental results have further improved, yielding an AUC score of 0.680 and DP_20 of 0.939. As for HOTC-2024-RedNIR, the performance of the baseline model “B” yields 0.328 and 0.428. Then, our proposed backbone network increased AUC and DP by 0.3% and 1.2%. It is worth noting that SAAM has further increased AUC and DP to 0.357 and 0.463, respectively. It demonstrates the effectiveness of $S^{2}$ in extracting spectral information, the superiority of SAFM over intra-modality attention, and the ability of SAAM to supplement the shortcomings of existing tracking heads in object localization prediction.
@@
TABLE III: ablation experimental results on the HOTC-2020 with the baseline model. $\Delta$ indicates the amount of variation.
@@
TABLE IV: ablation experimental results on the HOTC-2024-RedNIR with the baseline model. $\Delta$ indicates the amount of variation.
Methods AUC DP_20 $\Delta$(AUC) $\Delta$(DP_20)
B 0.328 0.428 - -
S 0.331 0.440 +0.003 +0.012
S+F 0.338 0.448 +0.010 +0.020
S+F+A 0.357 0.463 +0.029 +0.035
TABLE V: ablation experimental results of different backbone settings. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
TABLE VI: ablation experimental results of different fusion methods. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
TABLE VII: ablation experimental results of different prediction strategies. $\Delta$ indicates the amount of variation compared with the results of the first line.
@@
Evaluation of Different Backbone Settings. To assess the impact of varying backbone architectures, we conduct experiments with the individual spatial and spectral branches of the $S^{2}$ and remove the residual components. The experimental results are listed in Table V. We first verify the effectiveness of the residual structure, and the experimental results achieve 0.673 of the AUC score and 0.932 of DP_20, which shows the importance of the residual structure. Next, we explore the branches independently for feature extraction. The spatial branch, solely based on depthwise separable convolutions, lacks comprehensive channel information, resulting in a lower AUC score and DP_20 of 0.470 and 0.786, respectively. Meanwhile, the spectral branch alone achieved an AUC score of 0.658 and a DP_20 of 0.913, illustrating that while effective, the performance of each branch is not as robust as when combined. When conducting experiments with the complete $S^{2}$ structure, the model performance achieved an AUC score of 0.675 and DP_20 of 0.934, respectively. It demonstrates that the integrated approach of $S^{2}$ can capture both spatial and spectral information and offer a more comprehensive feature representation.
@@
Evaluation of Different Fusion Methods. The feature fusion is a key aspect of cross-modality tracking. To show the effectiveness of our proposed SAFM, we conduct experiments with various fusion methods, including element-wise addition (ADD), concatenation (CONCAT), and state-of-the-art spectral attention methods, including SE-Net [58], SaE-Net [59], and ECA-Net [43]. As presented in Table VI, the simplest methods, addition and concatenation, without any attention mechanisms, exhibit the least impressive performance. The current spectral attention mechanisms focus on enhancing within-modality features and fall short in harnessing cross-modality feature interactions. Consequently, the most effective result is offered by ECA-Net, yielding the AUC score and DP_20 of 0.673 and 0.932, respectively. However, these scores are still 0.2% lower than those achieved by SAFM. Such a comparison shows the advantages of SAFM in the context of feature fusion, confirming that our module not only refines intra-modality representation but also effectively captures and leverages the correlations across modalities.
@@
Evaluation of Different Prediction Strategies. In this section, we evaluate the performance of our weighted prediction approach. Drawing upon the “S” model configuration, we compare four different prediction strategies: using the HS branch alone, using the RGB branch alone, averaging the predictions from both, and applying a weighted prediction method. From Table VII, it can be seen that using the HS branch to predict only achieves an AUC score of 0.230 and DP_20 of 0.510. This was expected, as the $S^{2}$ is initially trained from scratch on the HOTC-2020 dataset, making it challenging for the HS branch alone to develop robust feature extraction capabilities. Then, the result of the RGB branch is 0.623 of the AUC score and DP_20 of 0.871. Due to the parameter freezing of the RGB branch, only the knowledge of the RGB domain is used for false-color image tracking, resulting in poor tracking performance. For average prediction, although the prediction results of both modalities are comprehensively considered, the tracking performance reached an AUC score of 0.639 and DP_20 of 0.915. However, in different scenarios, the contributions of HS and RGB modality to the results should be different. So, we used the weighted prediction method to adjust the weights of the two modality predictions adaptively, and the final results are the AUC score of 0.680 and DP_20 of 0.939. This further shows the effectiveness of the weighted prediction method.
Evaluation of Different Loss Weights. To explain the impact of the weights of the SAAL on the network’s performance, we conduct experiments with various weights $\gamma$ of SAAL, including 0.1, 0.5, 1.0, and 1.5, under the settings of $\alpha=1$ and $\beta=2$. The experimental results listed in Table VIII indicate that setting the weights $\gamma$ to 1.0 yields the best tracking performance. Hence, $\gamma$ is set to 1.0 in our experiments.
TABLE VIII: ablation experimental results of different weights of SAAL. $\Delta$ indicates the amount of variation compared with the results of the first line.
Methods AUC DP_20 $\Delta$(AUC) $\Delta$(DP_20)
0.1 0.330 0.431 - -
0.5 0.325 0.433 -0.005 +0.002
1.0 0.357 0.463 +0.027 +0.032
1.5 0.338 0.464 +0.008 +0.033